# AniMaker: Automated Multi-Agent Animated Storytelling with MCTS-Driven Clip Generation

### v1

@@ 4.4. Human Rating
Following MovieAgent’s evaluation framework (Wu et al., 2025), we conduct human evaluation with 10 participants on 90 storytelling videos from 9 models across 10 stories. Each video was rated on a 1-5 scale across five dimensions: Visual Appeal, Script Faithfulness, Narrative Coherence, Character Consistency, and Physical Law Adherence. Note that evaluating complete storytelling videos (rather than individual clips) typically yields lower scores due to the increased complexity and length. As shown in Table 5, our model achieves superior performance across all metrics, with particularly strong results in Character Consistency.
Table 5. Human rating results, presenting scores for Character Consistency (C.C.), Narrative Coherence (N.C.), Physical Law Adherence (P.L.), Script Faithfulness (S.F.), and Visual Appeal (V.A.) using a 1-5 scale.
Model C.C.$\uparrow$ N.C.$\uparrow$ P.L.$\uparrow$ S.F.$\uparrow$ V.A.$\uparrow$ Avg.$\uparrow$
@@ 4.5. Ablation Studies
We conduct two key ablation experiments. First, we ablate MCTS-Gen by setting $w_{1}=1,w_{2}=1$, essentially generating only one candidate per clip. This change results in a 7.1% reduction in performance (the green triangle in Figure 4) on AniEval, confirming the importance of our MCTS-driven generation strategy. Notably, even this ablated version still outperforms the best-performing baseline method by 6.6%, further demonstrating that our multi-agent framework remains highly competitive even without clip candidate selection. Next, we ablate AniEval by generating five candidates per clip with VBench for selection. This yields a score of 73.18, a 4.6% decrease compared to our full method. Qualitative assessment of the resulting videos also reveals noticeable degradation in action expressiveness and cross-clip consistency, highlighting the importance of AniEval for storytelling animation.
@@ 5. Limitations
Despite employing state-of-the-art models for storyboard and video clip generation, a gap remains between current model capabilities and commercial film production quality. Key limitations include poor adherence to physical laws and unrealistic interactions between scene elements, creating bottlenecks in production-ready output. However, our framework adopts a modular, plug-and-play architecture that enables seamless model integration. As more advanced models emerge, we will continuously update our framework and share these workflow improvements with the community, especially mitigating the hallucination of models (Wang et al., 2024).

### v2

@@ 4.4. Human Rating
Following MovieAgent’s evaluation framework (Wu et al., 2025), we conduct human evaluation with 10 participants on 90 storytelling videos from 9 models across 10 stories. Each video was rated on a 1-5 scale across five dimensions: Visual Appeal, Script Faithfulness, Narrative Coherence, Character Consistency, and Physical Law Adherence. Note that evaluating complete storytelling videos (rather than individual clips) typically yields lower scores due to the increased complexity and length. As shown in Table 5, our model achieves superior performance across all metrics, especially in terms of Character Consistency.
Table 5. Human rating results on a 1–5 scale, covering Character Consistency ($C.$), Narrative Coherence (N.C.), Physical-Law Adherence (P.L.), Script Faithfulness (S.F.), and Visual Appeal (V.A.).
[…]
@@ 4.5. Ablation Studies
We conduct two key ablation experiments. First, we ablate MCTS-Gen by setting $w_{1}=1,w_{2}=1$, essentially generating only one candidate per clip. This change results in a 7.1% reduction in performance (the green triangle in Figure 4) on AniEval, confirming the importance of our MCTS-driven generation strategy. Notably, even this ablated version still outperforms the best-performing baseline method by 6.6%, further demonstrating that our multi-agent framework remains highly competitive even without clip candidate selection. Next, we ablate AniEval by generating five candidates per clip with VBench for selection. This yields a score of 73.18, a 4.6% decrease compared to our raw method. Qualitative assessment of the resulting videos also reveals noticeable degradation in action expressiveness and cross-clip consistency, highlighting the importance of AniEval for storytelling animation.
@@ 5. Limitations
Despite employing state-of-the-art models for storyboard and video clip generation, a significant gap remains between current model capabilities and the quality required for commercial film production. A primary limitation is the restricted narrative scope and complexity our method can handle. Compared to the intricate cinematic grammar and narrative structures utilized in professional filmmaking, as detailed in (Ronfard, 2021), our approach currently falls short.
These shortcomings are rooted in several underlying defects within the video generation models. They are prone to producing animation artifacts, along with issues of incoherence and hallucination that disrupt narrative continuity. Furthermore, the models exhibit a relatively poor adherence to fundamental physical laws, resulting in unrealistic interactions between characters and scene elements, as also highlighted in (Shi et al., 2025). Luckily, our framework is built on a modular, plug-and-play architecture that allows for seamless model integration. As more advanced generative models emerge, we are committed to continuously updating our framework and sharing these workflow improvements with the community to address these limitations.
@@ 6. Conclusion
Acknowledgements.
We thank the editor and reviewers for their efforts in improving our paper. This work was supported by grants: Natural Science Foundation of China (No. 62422603), Guangdong Basic and Applied Basic Research and Foundation (No.2024B0101050003) and Shenzhen Science and Technology Progam (No. ZDSYS20230626091203008).
Appendix A LMM Prompts
This section presents the comprehensive set of prompts engineered for the LMM, which steers the entire automated animation creation pipeline. The figures below illustrate the sequence of these prompts, starting from the initial transformation of a narrative story into a structured script (Figure 7) and its corresponding validation (Figure 8). Following this, we detail the prompts for generating scene imagery (Figure 9), the video itself (Figure 10), and an analysis of the video generation prompt (Figure 11). The final set of prompts covers the creation (Figure 12), modification (Figure 13), and verification (Figure 14) of the voiceover script.
Appendix B Details of AniEval
AniEval enhances the evaluation framework of EvalCrafter (Liu et al., 2024a) through two main aspects. First, it automates the assessment of action related metrics by comparing user prompts with actions identified in the video clips. Second, it integrates several new metrics, including DreamSim (Fu et al., 2023), Count-Score (Cheng et al., 2024), and Face Consistency. For a comprehensive assessment, AniEval is organized into 4 primary domains comprising 14 fine-grained metrics, as detailed below.
B.1. Overall Video Quality
B.1.1. VQA_A and VQA_T
Dover (Wu et al., 2023) is employed to assess generated video quality, providing both an aesthetic score (VQA_A) and a technical score (VQA_T). The aesthetic score focuses on the overall visual appeal of the video, assessing aspects such as composition, color harmony, and artistic quality, while technical score evaluates common distortions like noise and artifacts.
B.1.2. MusIQ
MUSIQ (Ke et al., 2021) is adopted to evaluate the perceptual quality of the generated videos. Each clip is first sampled into a set of frames, after which the MUSIQ score is computed for every frame and averaged across the entire sequence. This yields a unified quality index that simultaneously captures global composition, local sharpness, and typical distortions such as blur and compression artifacts.
B.2. Text-Video Alignment
B.2.1. Text-Video Consistency
CLIPScore (Radford et al., 2021b) is employed to measure the semantic consistency between the input text prompt and the generated video. The pretrained ViT-B/32 CLIP model serves as the feature extractor, yielding text embeddings and frame-wise image embeddings; their cosine similarity is computed for every frame. The overall CLIPScore is obtained by averaging these per-frame similarities across the entire sequence.
B.2.2. Text-Story Consistency
To further evaluate the alignment between the input prompt and the generated video, a method leveraging BLIP2 (Li et al., 2023) and BLEU (Papineni et al., 2002) is also adopted. The BLIP2 model serves as a caption generator, producing five distinct text descriptions for each video. The BLEU score is then computed to measure the similarity between the original input prompt and each of these five generated captions. The overall score is the average of the five resulting BLEU scores.
B.2.3. Detection-Score
For each prompt, Gemini first identifies the key object and the desired quantity. To quantify its presence in the video, A Detection Score is computed by the SAMTrack model [cheng2023segment]. SAMTrack performs object detection on uniformly sampled frames; each frame returns a binary result (1 if the object is detected, 0 otherwise). The Detection Score is the mean of these binary values across all sampled frames.
B.2.4. Count-Score
Furthermore, to evaluate whether the video contains the correct number of objects, a Count Score is calculated. Using the same method, the number of detected objects in each frame is counted. This number is compared to the target quantity, yielding a binary frame-level score: 1 if the counts match, and 0 otherwise. The Count Score is the mean of these binary values across all sampled frames.
B.3. Video Consistency
B.3.1. DreamSim
DreamSim (Fu et al., 2023) is employed to measure the overall similarity between a given clip and its adjacent clips. For each comparison (i.e., between the current clip and its preceding or succeeding clip), frames are extracted from both. For every frame in the current clip, its minimum distance to any frame in the adjacent clip is calculated by Dreamsim. The final DreamSim score for the pair is the average of these distances.
B.3.2. Face Consistency
The Face Consistency metric evaluates animated character facial consistency using an InceptionNext (Yu et al., 2024) model trained on the Anime Face Dataset (splcher, 2019). The process begins by employing a Segment-and-Track-Anything (SAM) based approach to detect and extract character faces from video clips and reference image, guided by a text prompt. Each extracted face is then passed through the InceptionNext model to obtain a high-dimensional feature embedding, where the L2 distance between embeddings serves as a measure of dissimilarity. The metric performs two types of evaluation: for Reference-to-Video Consistency, it calculates the average L2 distance between the reference image’s face embedding and all face embeddings from the video frames. For Temporal Consistency between adjacent video clips featuring the same character, it determines the L2 distance from each face in the current clip to any face in the adjacent clip, and the final score is the average of these distances.
B.3.3. Warping Error
A pre-trained optical flow estimation network (Teed and Deng, 2020) is employed to obtain the optical flow between pairs of frames. Using this information, one frame is warped to align with the next. The pixel-wise difference between the warped image and the actual target frame is then calculated. Warp differences are computed for every frame pair, and the final score is obtained by averaging these results across the entire sequence.
B.3.4. Semantic Consistency
Beyond pixel-wise error, semantic consistency is also evaluated to assess the consistency between consecutive frames. Specifically, this involves calculating the cosine similarity between the feature embeddings (extracted by ViT-B/32 CLIP model) of each pair of adjacent frames. The final semantic consistency score is then derived by averaging these similarity values across all consecutive frame pairs in the video.
B.4. Motion Quality
For every prompt, Gemini is used to determine:
mm_action: Describe the primary action of the main character – e.g., “walking”, “talking”, “hugging”.
raft_amp: The perceived motion speed – “fast” or “slow”.
B.4.1. Action Recognition
The evaluation process begins by using the MMAction2 (Contributors, 2020) toolbox and the VideoMAE V2 model (Wang et al., 2023a) to identify the top five most probable actions and their corresponding confidence scores. Subsequently, a CLIP model calculates the textual similarity between these five predictions and the target action (mm_action). The final score is a weighted sum of these similarities, with the confidence scores serving as the weights.
B.4.2. Action Strength
The general action strength of the generated video is considered to ensure it is not overly static. To achieve this, the RAFT model (Teed and Deng, 2020) is first employed to extract the dense optical flows between each pair of frames. An average flow score is then computed by averaging these dense flows across the entire video clip.
B.4.3. Motion AC-Score
The alignment between a video’s motion and the prompt is evaluated using the average optical-flow score. A classification of ”fast” is assigned if the score exceeds the empirical threshold of $\rho$ = 5; otherwise, it is labeled ”slow.” A final comparison is then made between this label and the motion level required by the prompt.
Figure 7. Prompt for Story to Script Generation
Figure 8. Prompt for Story to Script Check
Figure 9. Prompt for Scene Image Generation
Figure 10. Prompt for Video Generation
Figure 11. Prompt for Video Generation Prompt Analysis
Figure 12. Prompt for Voiceover Script
Figure 13. Prompt for Voiceover Script Modification
Figure 14. Prompt for Voiceover Script Check